OpenStudy (anonymous):

Write each complex number in strandard form. √-5

OpenStudy (anonymous):

0 + (root5)i

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

yes i think so

OpenStudy (anonymous):

so i put o+√5

OpenStudy (anonymous):

0+√5

OpenStudy (anonymous):

√-1 *√-52 =i2√13

OpenStudy (anonymous):

i2√13 is the answer?

OpenStudy (anonymous):

sorry missing side √-1 *√52 = i2√13 yes

OpenStudy (anonymous):

i for negative side

OpenStudy (anonymous):

thank

OpenStudy (anonymous):

what??? i-

OpenStudy (anonymous):

when the root negative

OpenStudy (anonymous):

can you type how my answe is going to look

OpenStudy (anonymous):

√-52 =√-1=√-1 √-1= i =√52 = 2*2*13 = 2√13 than a answer is : 2i √13 or u can write : - 2√13 it mean negative put out side,keep radical square positive side

OpenStudy (anonymous):

okay so what about this one ? √-18

OpenStudy (anonymous):

sorry I keep tye wrong again √-52 =√-1*√52 √-1= i =√52 = 2*2*13 = 2√13 than a answer is : 2i √13 or u can write : - 2√13 it mean negative put out side,keep radical square positive sidea minute ago

OpenStudy (anonymous):

√-18= √-1*√18 = √-1= i = √18 = 2*3*3 = 3√2 = √-1 * 3√2 = 3i √2

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

2 way u can write, or wirte: -3√2 , keep in mined square root alway positive